# Profiling Bias in LLMs: Stereotype Dimensions in Contextual Word Embeddings

### v1

@@ 1 Introduction
Figure 1: 7D stereotype profile for Llama-3-8B, revealing differences in embeddings of 100 female and 100 male associated names. *Bold dimensions are statistically significant (p<0.05).
@@
Table 1: Examples of terms and their directions on stereotype dimensions from the theoretically grounded dictionary by Nicolas et al. (2021). The additional terms were collected from their extended dictionary created by a semi-automated method. Warmth: Sociability & Morality. Competence: Ability & Agency.
@@
Past research on bias has often involved word embedding association tests (Caliskan et al., 2017), inspired by the implicit association tests (IAT) (Greenwald et al., 1998) of social psychology. By another inspiration from the social sciences, a newer direction of natural language processing (NLP) research transforms in-transparent embeddings into a space of meaningful dimensions (Mathew et al., 2020; Kwak et al., 2021; Şenel et al., 2022; Engler et al., 2022), enabling new ways to study concepts. Similar to semantic differentials (Osgood et al., 1957), this methodology relies on antonyms or opposing concepts described by lexicons.
In this work we study bias in LLMs by transforming their embeddings based on the stereotype content model (SCM) by Fiske et al. (2002), enabling the study along theoretically and empirically grounded stereotype dimensions. The SCM entails two primary dimensions which originate from interactions, where people seek to understand the other party’s intent (dimension of warmth) and their capabilities (dimension of competence). Stereotypically women are thereby associated with higher warmth, men with higher competence. Furthermore we employ the extended model (Abele et al., 2016; Ellemers, 2017; Goodwin, 2015; Koch et al., 2016) allowing us to provide detailed 7D bias profiles as shown in Figure 1.
@@
In summary, we (i) show how the stereotype content model can be employed to expose and visualize bias in contextual embeddings, (ii) generate bias profiles for twelve LLMs for gender-associated names and gendered terms, displaying overall stereotypical associations of warmth and competence, (iii) provide insights on stereotype dimensions and bias across context examples and network layers.
@@ 2 Related Work
Inspired by the human implicit association test (Greenwald et al., 1998), Caliskan et al. (2017) developed the first Word Embedding Association Test (WEAT) to assess the association between two target concepts (e.g. scientist vs. librarian) and two attributes (e.g. male vs. female) in static word embeddings by cosine similarity and a permutation test. Later Tan and Celis (2019) built a first approach to measure bias for LLMs using the contextual embeddings of the words within examples. The Contextual Embedding Association Test (CEAT) (Guo and Caliskan, 2021) employs a random effects model to quantify bias with sampled contexts from a corpus.
@@
A newer approach to interpreting the high-dimensional embedding spaces is again inspired by a concept from the social sciences; semantic differentials (Osgood et al., 1957). Mathew et al. (2020) introduced POLAR, a transformation of static word embeddings to a new polar, interpretable space. The polar opposites are antonyms such as hot-cold or soft-hard and their word vectors are employed to define the new dimensions, which were shown to align with human judgement in an evaluation study. Similar frameworks are SemAxis (An et al., 2018), FrameAxis (Kwak et al., 2021) and BiImp (Şenel et al., 2022).
@@
More recently, the SensePolar framework was introduced by Engler et al. (2022), extending the POLAR approach to contextual word embeddings. The poles are hereby defined not by the word alone but by their embedding within sense-specific example sentences from a dictionary. The authors showed that these more interpretable embeddings can achieve similar performance to regular ones on natural language understanding (NLU) tasks, and furthermore evaluated their approach by a human evaluation study.
@@
Most similar to our work Fraser et al. (2021) analyze stereotype dimensions in static embeddings, combining the POLAR framework by Mathew et al. (2020) with the warmth and competence dimensions of the stereotype content model (Fiske et al., 2002). They demonstrate that static word embeddings can recreate the stereotype dimensions from literature by predicting the cold-warm and competent-incompetent associations for additional known words, and by further comparing the results to psychological surveys.
@@
For contextual embeddings Ungless et al. (2022) measure bias with CEAT (Guo and Caliskan, 2021) based on the warmth and competence dimensions and in a generation based approach Jeoung et al. (2023) elicit evaluation of different social groups on these dimensions, with multiple prompting strategies. The stereotype content model has also be used for de-biasing methods (Ungless et al., 2022; Omrani et al., 2023). The researchers suggest that this theory-driven approach has an advantage, because it is social-group-agnostic and thus does not require iteration over discriminated groups or previous knowledge of specific bias.
@@
In another projection approach, Omrani Sabbaghi et al. (2023) use a maximum margin support vector classifier to learn the valence subspace (pleasantness vs unpleasantness) and project the word ‘person’ to this dimension, placing different words in its context. The bias between words is then measured by their effect on the contextualized representation of ‘person’.
@@ 3.1 Stereotype Dimensions & Dictionaries
Our analysis of stereotype dimensions and bias in LLMs is grounded in the stereotype content model by Fiske et al. (2002), who showed that there are two major dimensions of warmth and competence, and that many stereotypes are mixed along these two. Following Nicolas et al. (2021) we also study the more fine-grained dimensions of sociability and morality for warmth (Abele et al., 2016) and ability and agency for competence (Ellemers, 2017; Goodwin, 2015). Expanding the set of concepts by the Agency-Beliefs-Communion model (Koch et al., 2016) we further include the dimensions of status, politics and religion. This allows us to also provide a more detailed and extended stereotype profile for the LLMs.
@@
The “seed dictionary” is theory-driven, using terms from literature, while the “full dictionary” contains additional terms collected by a semi-automated method, identifying synonyms using the English lexical database WordNet (Miller et al., 1990) [^2] https://wordnet.princeton.edu/. Both dictionaries distinguish seven stereotype dimensions, as shown in Table 1 with examples of terms per dimension and direction. While most dimensions are coded low–high, the politics dimension is coded progressive–traditional and the dimension of religion is coded non-religious–religious.
Figure 2: Properties of context examples.
@@ 3.2 Context Examples
Generated Examples: For our main experiments, we generate gender-neutral contexts with Llama-3-8B-instruct by the instruction to avoid names and gender-specific pronouns. Thus, no additional gender bias is introduced. As synset information is available for the terms in the seed dictionary, the prompts for these terms additionally include the term definition from WordNet (Miller et al., 1990), allowing a more precise generation for the specific word meaning.
Dictionary Examples: As in the original SensePolar paper (Engler et al., 2022) we also retrieve context examples from WordNet (Miller et al., 1990), and only for the seed stereotype dictionary we manually add examples from other dictionaries where WordNet does not provide any.
@@
Example properties are shown in Figure 2. We fixed the number of examples to five for comparison and to limit computational time, but for the dictionary examples, there often are fewer available. Dictionary examples are also the shortest, as they are often short phrases, e.g. “friendly advice”. The generated examples are the longest with an average of 20 words. Reddit examples are truncated on both sides, to include context around the term, which may explain later term positions.
@@
Llama-3-8B (AI@Meta, 2024) 0.79 0.81 0.66 0.87 0.72 0.77 0.8 0.73 0.74
Llama-3-8B-Instruct (AI@Meta, 2024) 0.79 0.82 0.65 0.88 0.75 0.77 0.83 0.83 0.81
Llama-3.2-3B (AI@Meta, 2024) 0.81 0.8 0.63 0.87 0.66 0.76 0.79 0.77 0.62
Llama-3.2-3B-Instruct (AI@Meta, 2024) 0.78 0.84 0.62 0.88 0.75 0.78 0.81 0.86 0.62
Gemma-2B (TeamGemma et al., 2024) 0.66 0.67 0.58 0.75 0.66 0.69 0.78 0.58 0.96
Gemma-2-2B (TeamGemma et al., 2024) 0.7 0.67 0.64 0.8 0.74 0.7 0.75 0.63 0.95
@@
Table 2: Accuracy for the direction prediction task. Hereby, we predict the direction on the stereotype dimensions (e.g. high warmth vs. low warmth) for terms in the extended stereotype dictionary (Nicolas et al., 2021) and compare to ground truth directions.
@@ 3.3 Polar Projection
For the computation of stereotype dimensions we follow the SensePolar framework by Engler et al. (2022) which is an extension of the POLAR framework (Mathew et al., 2020) for contextual embeddings. Hereby the embeddings are transformed into an interpretable space based on polar dimensions, which in our case are defined by the stereotype dictionary. We transform the embeddings at two levels: (i) Warmth + competence, and (ii) seven granular dimensions of the extended stereotype content model.
Similar to Fraser et al. (2021) we take the words for each stereotype dimension from the theory-driven “seed dictionary” (Nicolas et al., 2021) and we average individually the word embeddings for the high and for the low classified words, for which the numbers are shown in Table 1. Word lists for warmth (sociability + morality) and competence (ability + agency) are compiled of the words of their subordinate dimensions.
As a first step we calculate the sense embedding $\mathbf{s}$ for a word with its specific sense and m sense-specific context examples, as shown in Equation 1. We hereby average the contextual embeddings $\mathbf{w}$ across the different context examples, also averaging across subwords when words are split due to subword tokenization.
@@
For n words belonging to the pole of a stereotype dimension, e.g. “friendlines” and “sociability” for the pole “high sociability”, we average their sense embeddings to obtain the average pole embedding $\mathbf{p}$. Next, we stack the vectors and subtract the low dimension embeddings from the high dimension embeddings to obtain the change of basis matrix $\mathbf{a}$, describing the newly defined space with $\mathbf{h}$ stereotype dimensions:
@@
The new embedding $\mathbf{d}$ in the stereotype space can be interpreted as follows: Similar to the semantic differential technique, a higher absolute value signifies a stronger semantic relation to the respective dimension. A positive value signifies an association with the high pole, for example “high morality” and a negative value signifies an association with the low pole, e.g. “low morality”.
@@
To evaluate the consistency of the stereotype dimensions we follow the approach by Fraser et al. (2021) and project additional terms from the full dictionary by Nicolas et al. (2021) to the stereotype space. For each term we predict its direction on its assigned dimension by the sign of its polar value and calculate the accuracy against the ground truth directions of the words.
@@
For the analysis of gender bias we project gender-associated words to our stereotype dimensions, utilizing two larger binary ‘vocabulary populations’ and individual terms for transgender and nonbinary gender. The largest population are historically gender-associated names such as “Mary” and “John”, taken from the most popular given names of the last century in the United States [^4] https://www.ssa.gov/oact/babynames/decades/century.html; 100 for each gender. Then we employ gendered terms by definition, e.g. “woman”, “she” and “man”, “he”, as utilized in experiments of WEAT (Caliskan et al., 2017). As context for our gendered terms and names, we use neutral templates, such as “This is [NAME]” or “This is [TERM]” to provide context without unnecessarily introducing additional bias (compare May et al. (2019); Tan and Celis (2019)). We average across representations for the different templates to achieve a more robust representation for the names and terms.
For easier interpretation and comparison between models we standardize the projected polar values, separately for names and terms, as preliminary work showed that named entities and pronouns can show different average tendencies on stereotype dimensions. To assesses the significance of the observed differences we employ t-tests.
@@ 3.4 Models
For our analysis we project open source models of multiple generations available in the Huggingface Library [^5] https://huggingface.co/ onto stereotype dimensions. Model details are included in Table 2.
@@ 4 Results
Figure 3: 2D Stereotype Profiles. LW/HW = Low/High Warmth. LC/HC = Low/High Competence. *Significant dimensions (p<0.05).
Figure 4: 2D Stereotype profile for Llama-3-8B with nonbinary terms. LW/HW = Low/High Warmth. LC/HC = Low/High Competence.
@@ 4.1 Prediction of Direction for Additional Terms
Predicting the direction on the stereotype dimensions for new terms from the extended stereotype dictionary (Nicolas et al., 2021), we find most studied models can perform this task well, with different strengths. Table 2 shows the results when embedding and projecting with the generated examples. OlMo-1B-hf (warmth and competence) and Llama3.2-3B-Instruct (only warmth) achieve the performance closest to that of static embeddings by Fraser et al. (2021), where the FastText-based model scored respectively 0.85 for warmth and 0.86 for competence. OlMO and the Llama models also perform very well for the granular dimensions. Predicting morality is the overall easiest task for the models, while the other subdimension of warmth, sociability, poses the most difficult task.
Only GPT-Neo-125M performs worse than chance for some dimensions, however, this pertains only to raw polar values. When we use a different cut-off by mean-centering the values, the model achieves much better results, e.g. 0.76 and 0.71 for warmth and competence. Similarly, GPT2 and the Gemma models benefit from a different cut-off value, gaining up to 10 percentage points per dimension. Thus all models are able to discriminate between low and high-labeled terms on the stereotype dimensions.
Figure 5: 7D stereotype profiles for gender-associated names and gendered terms with Llama-3.2-3B-instruct. *Bold dimensions are statistically significant (p<0.05).

### v2

@@ 1 Introduction
Figure 1: 7D stereotype profile for Llama-3-8B, revealing differences in embeddings of 100 female and 100 male-associated names. *Statistically significant differences (p<0.05).
@@
Table 1: Examples of terms and their directions on stereotype dimensions from the theoretically grounded dictionary by Nicolas et al. (2021). The additional terms were collected from their extended dictionary created by a semi-automated method. High-level stereotype dimensions are constructed as follows: warmth = sociability + morality, competence = ability + agency.
@@
Past research on bias has often involved word embedding association tests (Caliskan et al., 2017), inspired by the implicit association tests (IAT) (Greenwald et al., 1998) of social psychology. By another inspiration from the social sciences, a newer direction of natural language processing (NLP) research transforms opaque embeddings into a space of meaningful dimensions (Mathew et al., 2020; Kwak et al., 2021; Şenel et al., 2022; Engler et al., 2022), enabling new ways to study concepts. Similar to semantic differentials (Osgood et al., 1957), this methodology relies on antonyms (e.g. fast vs. slow) or opposing concepts described by lexicons.
In this work we study bias in LLMs by transforming their embeddings based on the stereotype content model (SCM) by Fiske et al. (2002), enabling the study along theoretically and empirically grounded stereotype dimensions. The SCM entails two primary dimensions originating from interactions, where people seek to understand the other party’s intent (dimension of warmth) and their capabilities (dimension of competence). Stereotypically, women are thereby associated with higher warmth, and men with higher competence. Furthermore we employ the extended model (Abele et al., 2016; Ellemers, 2017; Goodwin, 2015; Koch et al., 2016), allowing us to provide detailed 7D bias profiles as shown in Figure 1.
@@
In summary, we (i) show how the stereotype content model can be employed to expose and visualize bias in contextual embeddings [^1] Code available at https://github.com/carolinmschuster/profiling-bias-in-llms, (ii) generate bias profiles for twelve LLMs for gender-associated names and gendered terms, displaying overall stereotypical associations of warmth and competence, (iii) provide insights on stereotype dimensions and gender bias across context examples and network layers.
@@ 2 Related Work
Inspired by the human implicit association test (Greenwald et al., 1998), Caliskan et al. (2017) developed the first Word Embedding Association Test (WEAT) to assess the association between two target concepts (e.g., scientist vs. librarian) and two attributes (e.g., male vs. female) in static word embeddings by cosine similarity and a permutation test. Later Tan and Celis (2019) built a first approach to measure bias for LLMs using the contextual embeddings of the words within examples. The Contextual Embedding Association Test (CEAT) (Guo and Caliskan, 2021) employs a random effects model to quantify bias with sampled contexts from a corpus.
@@
A newer approach to interpreting the high-dimensional embedding spaces is again inspired by a concept from the social sciences; semantic differentials (Osgood et al., 1957). Mathew et al. (2020) introduced POLAR, a transformation of static word embeddings to a new polar, interpretable space. The polar opposites are antonyms such as hot-cold or soft-hard, and their word vectors are employed to define the new dimensions, which were shown to align with human judgment in an evaluation study. Similar frameworks are SemAxis (An et al., 2018), FrameAxis (Kwak et al., 2021) and BiImp (Şenel et al., 2022).
@@
More recently, the SensePolar framework was introduced by Engler et al. (2022), extending the POLAR approach to contextual word embeddings. The poles are hereby defined not by the word alone but by their embedding within sense-specific example sentences from a dictionary. The authors showed that these more interpretable embeddings can achieve similar performance to regular ones on natural language understanding (NLU) tasks, and furthermore confirmed the approach by a human evaluation study.
@@
Most similar to our work Fraser et al. (2021) analyzed stereotype dimensions in static embeddings, combining the POLAR framework by Mathew et al. (2020) with the warmth and competence dimensions of the stereotype content model (Fiske et al., 2002). They demonstrated that static word embeddings can recreate the stereotype dimensions from literature by predicting the cold-warm and competent-incompetent associations for additional known words, and by further comparing the results to psychological surveys.
@@
For contextual embeddings Ungless et al. (2022) measured bias with CEAT (Guo and Caliskan, 2021) based on the warmth and competence dimensions and in a generation-based approach Jeoung et al. (2023) elicited evaluation of different social groups on these dimensions, with multiple prompting strategies. The stereotype content model has also been used for de-biasing methods (Ungless et al., 2022; Omrani et al., 2023). The researchers suggest that this theory-driven approach has an advantage because it is social-group-agnostic and thus does not require iteration over discriminated groups or previous knowledge of specific bias.
@@
In another projection approach, Omrani Sabbaghi et al. (2023) used a maximum margin support vector classifier to learn the valence subspace (pleasantness vs. unpleasantness) and projected the word ‘person’ to this dimension, placing different words in its context. The bias between words is measured by their effect on the contextualized representation of ‘person’.
This work furthermore relates with a broader range of studies trying to understand the contents of contextual representations, most notably by knowledge probing (e.g. Tenney et al. (2019b); Schuster and Hegelich (2022)). See Cao et al. (2024) for a recent survey.
@@ 3.1 Stereotype Dimensions & Dictionaries
Our analysis of stereotype dimensions and bias in LLMs is grounded in the stereotype content model by Fiske et al. (2002), who showed that there are two major dimensions of warmth and competence and that many stereotypes are mixed along these two. Following Nicolas et al. (2021) we also study the more fine-grained dimensions of sociability and morality for warmth (Abele et al., 2016) and ability and agency for competence (Ellemers, 2017; Goodwin, 2015). Expanding the set of concepts by the Agency-Beliefs-Communion model (Koch et al., 2016), we further include the dimensions of status, politics, and religion. This allows us to also provide a more detailed and extended stereotype profile for the LLMs.
@@
The ‘seed dictionary’ is theory-driven, using terms from literature, while the ‘full dictionary’ contains additional terms collected by a semi-automated method, identifying synonyms using the English lexical database WordNet (Miller et al., 1990) [^3] https://wordnet.princeton.edu/. Both dictionaries distinguish seven stereotype dimensions, as shown in Table 1 with examples of terms per dimension and direction. While most dimensions are coded low–high, the politics dimension is coded progressive–traditional and the dimension of religion is coded non-religious–religious.
Figure 2: Properties of context examples: Histograms of example counts, numbers of words and positions of dictionary terms within the examples.
@@ 3.2 Context Examples
Generated Examples: For our main experiments, we generate gender non-specific contexts with Llama-3-8B-instruct AI@Meta (2024a); Dubey et al. (2024) by the instruction to avoid names and gender-specific pronouns. Thus, no additional gender bias is introduced. As synset information is available for the terms in the seed dictionary, the prompts for these terms additionally include the term definition from WordNet (Miller et al., 1990), allowing a more precise generation for the specific word meaning.
Dictionary Examples: As in the original SensePolar paper (Engler et al., 2022), we also retrieve context examples from WordNet (Miller et al., 1990), and only for the seed stereotype dictionary do we manually add examples from other dictionaries where WordNet does not provide any.
@@
Example properties are shown in Figure 2. We set the number of examples to five for comparison and to limit computational time, but there often are fewer available for the dictionary examples. Dictionary examples are also the shortest, as they are often short phrases, e.g. “friendly advice”. The generated examples are the longest with an average of 20 words. Reddit examples are truncated on both sides to include context around the term, which may explain later term positions.
@@
Llama-3-8B (AI@Meta, 2024a) 0.79 0.81 0.66 0.87 0.72 0.77 0.8 0.73 0.74
Llama-3-8B-Instruct (AI@Meta, 2024a) 0.79 0.82 0.65 0.88 0.75 0.77 0.83 0.83 0.81
Llama-3.2-3B (AI@Meta, 2024b) 0.81 0.8 0.63 0.87 0.66 0.76 0.79 0.77 0.62
Llama-3.2-3B-Instruct (AI@Meta, 2024b) 0.78 0.84 0.62 0.88 0.75 0.78 0.81 0.86 0.62
Gemma-2B (TeamGemma et al., 2024a) 0.66 0.67 0.58 0.75 0.66 0.69 0.78 0.58 0.96
Gemma-2-2B (TeamGemma et al., 2024b) 0.7 0.67 0.64 0.8 0.74 0.7 0.75 0.63 0.95
@@
Table 2: Accuracy for the direction prediction task. Additional terms in the extended stereotype dictionary (Nicolas et al., 2021) are embedded and projected to the stereotype dimensions. Projected positive/negative values predict high/low direction, e.g. a value of -0.3 for warmth is registered as low warmth. The highest accuracy for each dimension is shown in bold. Please refer to Table 1 for examples of terms with high and low labels for each dimension.
@@ 3.3 Polar Projection
For the computation of stereotype dimensions, we follow the SensePolar framework by Engler et al. (2022), which is an extension of the POLAR framework (Mathew et al., 2020) for contextual embeddings. Hereby, the embeddings are transformed into an interpretable space based on polar dimensions, which, in our case, are defined by the stereotype content dictionary. We transform the embeddings at two levels: (i) Warmth + competence, and (ii) seven granular dimensions of the extended stereotype content model.
Similar to Fraser et al. (2021), we take the words for each stereotype dimension from the theory-driven ‘seed dictionary’ (Nicolas et al., 2021), and we average individually the word embeddings for the high and for the low classified words, for which the numbers are shown in Table 1. Word lists for warmth (sociability + morality) and competence (ability + agency) are compiled of the words of their subordinate dimensions.
As a first step, we calculate the sense embedding $\mathbf{s}$ for a word with its specific sense and m sense-specific context examples, as shown in Equation 1. We hereby average the contextual embeddings $\mathbf{w}$ across the different context examples, also averaging across subwords when words are split due to subword tokenization.
@@
For n words belonging to the pole of a stereotype dimension, e.g. “friendliness” and “sociability” for the pole “high sociability”, we average their sense embeddings to obtain the average pole embedding $\mathbf{p}$. Next, we stack the vectors and subtract the low-dimension embeddings from the high-dimension embeddings to obtain the change of basis matrix $\mathbf{a}$, describing the newly defined space with $\mathbf{h}$ stereotype dimensions:
@@
Regarding the warmth and competence transformation, there are only two direction vectors. If, for example, the original contextual embedding has 768 dimensions, the change of basis matrix $\mathbf{a}$ has a shape of (2, 768).
@@
The new embedding $\mathbf{d}$ in the 2D or 7D stereotype space can be interpreted as follows: Similar to the semantic differential technique, a higher value signifies a higher association with the high pole, for example, “high morality”, and a lower value signifies a more significant association with the low pole, e.g., “low morality”. By projecting multiple terms we can compare their differences on these dimensions.
@@
To evaluate the consistency of the stereotype dimensions we follow the approach by Fraser et al. (2021) and project additional terms from the extended dictionary by Nicolas et al. (2021) to the stereotype space. Hereby, we use the same types of context examples as for the polar space creation. For each term, we predict its direction on its assigned dimension by the sign of its polar value, e.g., a value of -0.5 for sociability is registered as low sociability. To calculate the accuracy, we compare these predictions against the labels in the dictionary.
@@
For the analysis of gender bias, we project gender-associated words to our stereotype dimensions, utilizing two larger binary ‘vocabulary populations’ and individual terms for transgender and nonbinary gender (see Figure 4).
The largest populations are 100 historically female-associated names (e.g., Mary, Patricia) and 100 male-associated names (e.g., James, Michael), taken from the most popular given names of the last century in the United States [^5] https://www.ssa.gov/oact/babynames/decades/century.html.
Second, we employ binary gendered terms by definition as utilized in experiments of WEAT (Math vs. Arts and Science vs. Art) (Caliskan et al., 2017). For each gender, we project nine terms:
  • Female terms: female, woman, girl, sister, she, daughter, mother, aunt, grandmother
  • Male terms: male, man, boy, brother, he, son, father, uncle, grandfather
As examples for our gendered terms and names, we use neutral templates, such as “This is [NAME]” or “This is [TERM]” to provide context without unnecessarily introducing additional bias (compare May et al. (2019); Tan and Celis (2019)). We average across the different templates for a more robust contextual representation of names and terms.
For easier interpretation and comparison between models, we standardize the projected polar values separately for names and terms, as preliminary work showed that named entities and pronouns can show different average tendencies on stereotype dimensions. To assess the significance of the observed differences, we employ t-tests.
@@ 3.4 Models
For our evaluation, we project open source models of multiple generations available in the Huggingface Library [^6] https://huggingface.co/ onto stereotype dimensions. Model names and references are shown in Table 2. Except for the layer-wise analysis, we extract their average contextual representations across all layers, including the first embedding layer.
@@ 4.1 Prediction of Direction for Additional Terms
Predicting the direction on the stereotype dimensions for new terms from the extended stereotype dictionary (Nicolas et al., 2021), we find most studied models can perform this task well, with different strengths. Table 2 shows the results when embedding and projecting with the generated examples. OlMo-1B-hf (warmth and competence) and Llama3.2-3B-Instruct (only competence) achieve the performance closest to that of static embeddings by Fraser et al. (2021), where the FastText-based model scored respectively 0.85 for warmth and 0.86 for competence. OlMO and the Llama models also perform very well for the granular dimensions. Predicting morality is the overall easiest task for the models, while the other subdimension of warmth, sociability, poses the most difficult task. Accuracy varies greatly for religion, where there are 142 high, but only 6 low-labeled additional terms.
Only GPT-Neo-125M performs worse than chance for some dimensions, however, this pertains only to raw polar values. When we use a different cut-off than zero for predicting low/high directions by mean-centering the projected values, the model achieves much better results, e.g., 0.76 accuracy for warmth and 0.71 for competence. Similarly, GPT2 and the Gemma models benefit from a mean-based cut-off value, gaining up to 10 percentage points per dimension. Thus projections are spread on different ranges of values, but all models can reasonably discriminate between low and high-labeled terms on the stereotype dimensions.
Figure 3: 2D stereotype profiles for 100 female/male-associated names (left) and 9 female/male gendered terms (right).
LW/HW = Low/High Warmth. LC/HC = Low/High Competence. *Dimensions with statistically significant differences (p<0.05).